# Towards a clinically accessible radiology foundation model: open-access and lightweight, with automated evaluation

### v4

@@ Methods
Small Multimodal Model LLaVA-Rad follows LLaVA [31, 30] and LLaVA-Med [29] which leverage the capabilities of a pre-trained image encoder and a pre-trained language model to create a SMM. We choose BiomedCLIP-CXR as our image encoder and Vicuna-7B-v1.5 [9] as our language model. A multi-layer perceptron (MLP), which is randomly initialized, is introduced to project image features extracted by the image encoder into the word embedding space of the language model. Conditioned on the projected image features (visual tokens) and textual tokens, LLaVA-Rad generates text in an autoregressive manner. We refer the reader to LLaVA [31, 30] for a more in-depth description of the model architecture.
@@
Variant #1: Stage 1 pre-trains the image encoder on MIMIC-CXR only. Stage 2&3 are the same.
Variant #2: No stage-1 pretraining. Stage 2&3 are the same.
Variant #3: Stage 1 is the same. Stage 2&3 only use rule-processed MIMIC-CXR training data.
Variant #4: Stage 1 is the same. Stage 2&3 only use GPT-4 processed MIMIC-CXR training data.
Variant #5: Stage 1&2 are the same. No stage 3.

### v5

@@ Methods
Small Multimodal Model LLaVA-Rad leverages the capabilities of a pre-trained image encoder and a pre-trained language model to create a SMM. We choose BiomedCLIP-CXR as our image encoder and Vicuna-7B-v1.5 [9] as our language model. A multi-layer perceptron (MLP), which is randomly initialized, is introduced to project image features extracted by the image encoder into the word embedding space of the language model. Conditioned on the projected image features (visual tokens) and textual tokens, LLaVA-Rad generates text in an autoregressive manner. We refer the reader to LLaVA [31, 30] for a more in-depth description of the model architecture.
@@
Variant #1: Stage 1 pre-trains the image encoder on MIMIC-CXR only. Stage 2 and 3 are the same.
Variant #2: No stage-1 pretraining. Stage 2 and 3 are the same.
Variant #3: Stage 1 is the same. Stage 2 and 3 only use rule-processed MIMIC-CXR training data.
Variant #4: Stage 1 is the same. Stage 2 and 3 only use GPT-4 processed MIMIC-CXR training data.
Variant #5: Stage 1 and 2 are the same. No stage 3.